OpenStudy (anonymous):

the answer i got was 56 I just wanted to see if i was doing this right

OpenStudy (anonymous):

Then answer questions people ask when they try to help you.

OpenStudy (anonymous):

ah? that's wonderful! what might be that an answer to?

OpenStudy (saifoo.khan):

that wasnt the answer.

OpenStudy (bahrom7893):

what was the question?

OpenStudy (anonymous):

Probably this one.. 7/3*-4/3*(-6)

OpenStudy (anonymous):

not that ALK has stayed long enough to know for sure.

OpenStudy (bahrom7893):

so multiply out denominators and numerators, this is simple.

OpenStudy (anonymous):

Indeed.

OpenStudy (saifoo.khan):

i knw thats simple.

OpenStudy (saifoo.khan):

but still, my answer was deleted. :

OpenStudy (bahrom7893):

then i guess we should go someplace else.. I hate it when a user does that.. u spend at least 15 mins on a problem and the user just vanishes, no thank yous, no nothing..

OpenStudy (anonymous):

yep. That's also why I ask questions. If they're not going to answer a question in the first 2 mins, chances are they didn't stick around to work out the problem.

OpenStudy (anonymous):

so I can bail before I spend much time on it.

OpenStudy (curry):

polpak can you help me

OpenStudy (anonymous):

Certainly.. but only if you ask a question.

OpenStudy (curry):

well cause

OpenStudy (saifoo.khan):

are u referring this to me? so I can bail before I spend much time on it.

OpenStudy (curry):

its not just 1 question its a whole bunch

OpenStudy (curry):

its more like a online test that i need help

OpenStudy (saifoo.khan):

i can help too, unless my answer isnt deleted. :) + u got a cool picture!

OpenStudy (curry):

for 30 min striaght

OpenStudy (curry):

?

OpenStudy (anonymous):

Saif, if you don't want your answers deleted, there's a very simple fix. Don't give answers.

OpenStudy (anonymous):

Curry, I'm afraid that's not really what this site is for. I suggest studying for the test by doing practice problems and then see how well you do.

OpenStudy (curry):

so waht do you gusys say

OpenStudy (curry):

well its not realy a test aits practice thingi found onkine

OpenStudy (curry):

hey saifoo can you do it

OpenStudy (saifoo.khan):

lol, i can.. But im not allowed to give the answers.

OpenStudy (saifoo.khan):

u knw hindi right?

OpenStudy (curry):

i know just help me go to my page

OpenStudy (curry):

i know just help me go to my page

OpenStudy (curry):

i know just help me go to my page

OpenStudy (curry):

lthe last question i have asked

OpenStudy (curry):

click the last question then click the link on the bottom

OpenStudy (curry):

ok?

OpenStudy (saifoo.khan):

whenever you're dealing with e^x, you can use the natural log ln example: e^x = xln(e) bring the power down (here it's just x), and ln(e), which equals 1 e^(0.5x) = 0.5 0.5xln(e) = ln(0.5) 0.5x = ln(0.5) x = 2ln(0.5)

OpenStudy (curry):

you know let me send you the link ok click on it